# LimiX: Unleashing Structured-Data Modeling Capability for Generalist Intelligence

### v1

@@ Results.
OpenML-cc18 Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX 0.939±0.004 0.893±0.008 0.811±0.014 4.048 4.952 4.258 AutoGluon 0.932±0.000 0.885±0.001 0.790±0.003 5.823 6.694 7.048 TabPFN-v2 0.929±0.002 0.886±0.005 0.790±0.006 7.177 7.177 7.629 XGBoost 0.929±0.002 0.879±0.005 0.775±0.012 8.952 10.145 10.000 TabICL 0.927±0.004 0.875±0.008 0.782±0.011 8.242 9.677 9.613 LightGBM 0.927±0.003 0.879±0.005 0.775±0.015 9.323 9.532 9.565 CatBoost 0.926±0.005 0.870±0.008 0.770±0.016 11.258 12.726 12.613 RF 0.925±0.003 0.871±0.006 0.762±0.016 12.871 12.452 13.145 ET 0.922±0.003 0.861±0.006 0.721±0.017 12.903 15.016 16.839 Mitra 0.920±0.001 0.866±0.004 0.743±0.004 12.177 13.758 15.226 ExcelFormer 0.918±0.014 0.870±0.014 0.773±0.022 13.065 12.210 12.048 ResNet 0.913±0.009 0.860±0.011 0.764±0.019 13.435 13.613 12.806 TANGOS 0.910±0.007 0.863±0.010 0.759±0.025 13.355 13.339 13.484 MLP 0.908±0.007 0.857±0.008 0.743±0.019 13.887 13.984 15.016 T2G-Former 0.908±0.012 0.859±0.011 0.748±0.020 15.145 14.677 14.710 ModernNCA 0.906±0.009 0.858±0.013 0.747±0.021 16.661 16.129 15.532 FT-Transformer 0.904±0.015 0.856±0.012 0.739±0.019 16.532 15.306 16.194 TabR 0.900±0.015 0.863±0.011 0.757±0.025 16.661 13.484 13.113 DANets 0.900±0.021 0.840±0.029 0.715±0.053 18.145 17.613 18.581 DCN-v2 0.899±0.017 0.851±0.015 0.729±0.046 18.710 17.516 18.129 SNN 0.896±0.014 0.835±0.023 0.698±0.055 20.161 20.000 20.774 MLP-PLR 0.896±0.010 0.858±0.008 0.734±0.020 15.919 15.645 16.000 NODE 0.889±0.042 0.809±0.060 0.626±0.126 23.177 24.242 26.194 RealMLP 0.889±0.016 0.858±0.013 0.742±0.026 16.274 14.839 14.065 AutoInt 0.889±0.010 0.835±0.015 0.717±0.033 18.145 18.161 18.290 SwitchTab 0.881±0.031 0.819±0.039 0.671±0.065 21.403 20.790 21.516 TabCaps 0.875±0.024 0.848±0.017 0.692±0.042 21.597 18.613 20.161 TabNet 0.863±0.039 0.816±0.044 0.666±0.076 25.226 23.177 24.839 TabTransformer 0.829±0.021 0.776±0.022 0.616±0.051 22.565 21.645 22.387 GrowNet 0.816±0.027 0.748±0.027 0.581±0.050 25.403 25.855 24.935 SAINT 0.512±0.000 0.492±0.000 0.441±0.000 25.210 24.597 23.290
@@
Table 7: Classification results on the TabArena benchmark. The best scores are shown in bold.
TabArena Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX 0.849±0.005 0.877±0.004 0.597±0.014 3.182 3.030 6.606 AutoGluon 0.844±0.001 0.870±0.000 0.574±0.001 5.212 4.636 8.364 LightGBM 0.841±0.002 0.868±0.002 0.574±0.011 6.758 7.667 9.848 TabICL 0.840±0.006 0.870±0.002 0.553±0.007 6.364 5.727 9.970 XGBoost 0.838±0.002 0.867±0.003 0.567±0.018 7.576 6.939 10.182 TabPFN-v2 0.838±0.002 0.872±0.002 0.589±0.004 6.545 4.000 8.242 RF 0.837±0.002 0.864±0.003 0.558±0.011 9.000 7.455 11.303 CatBoost 0.835±0.005 0.867±0.004 0.574±0.015 9.333 6.758 9.273 ET 0.833±0.003 0.857±0.003 0.505±0.019 10.030 10.333 15.515 Mitra 0.815±0.025 0.862±0.008 0.533±0.049 11.152 9.485 14.152 ExcelFormer 0.810±0.014 0.849±0.015 0.555±0.027 13.788 15.939 14.152 TANGOS 0.791±0.019 0.844±0.009 0.522±0.047 17.636 16.848 15.212 TabR 0.785±0.012 0.842±0.013 0.510±0.049 19.182 18.606 15.303 ModernNCA 0.783±0.020 0.846±0.006 0.511±0.025 18.636 19.424 16.152 ResNet 0.781±0.029 0.824±0.034 0.532±0.045 19.394 20.515 15.515 MLP-PLR 0.781±0.007 0.836±0.004 0.460±0.013 18.121 18.091 18.818 T2G-Former 0.779±0.037 0.822±0.047 0.482±0.092 17.727 19.091 15.061 RealMLP 0.777±0.020 0.822±0.015 0.512±0.032 19.788 20.030 13.697 MLP 0.772±0.016 0.822±0.014 0.459±0.019 17.576 16.697 18.727 FT-Transformer 0.770±0.054 0.803±0.067 0.468±0.089 19.545 20.667 17.788 DCN-v2 0.769±0.037 0.833±0.019 0.482±0.051 20.333 18.424 17.394 AutoInt 0.769±0.042 0.826±0.032 0.474±0.060 18.788 19.636 17.788 NODE 0.769±0.073 0.792±0.070 0.352±0.145 19.333 20.515 23.485 SNN 0.755±0.054 0.818±0.033 0.442±0.054 21.273 21.091 21.152 SwitchTab 0.754±0.040 0.799±0.030 0.409±0.082 22.303 23.939 21.485 DANets 0.749±0.050 0.776±0.084 0.453±0.073 22.182 22.848 17.545 TabCaps 0.742±0.019 0.837±0.007 0.471±0.020 21.394 17.000 18.818 TabTransformer 0.739±0.021 0.781±0.022 0.438±0.040 19.667 20.424 18.303 TabNet 0.709±0.089 0.789±0.070 0.438±0.081 24.848 21.364 22.303 SAINT 0.694±0.021 0.739±0.018 0.437±0.029 23.576 24.061 20.606 GrowNet 0.646±0.052 0.674±0.083 0.361±0.072 25.697 27.091 22.394
@@
(a) AUC on the TabArena benchmark.
(b) Accuracy on the TabArena benchmark.
(c) F1-score on the TabArena benchmark.
Figure 11: Critical difference diagrams on the TabArena benchmark.
@@
PFN-CLS Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX 0.923±0.005 0.862±0.009 0.786±0.015 2.034 3.000 2.862 TabPFN-v2 0.910±0.002 0.845±0.007 0.756±0.010 4.862 6.793 6.483 AutoGluon 0.906±0.001 0.835±0.001 0.738±0.002 5.172 7.069 6.793 TabICL 0.903±0.005 0.832±0.008 0.742±0.014 6.276 8.034 8.379 XGBoost 0.898±0.003 0.831±0.007 0.733±0.023 8.241 6.862 7.759 Mitra 0.897±0.010 0.826±0.017 0.719±0.032 8.621 9.552 11.276 RF 0.896±0.006 0.822±0.008 0.721±0.019 10.414 10.172 11.310 CatBoost 0.895±0.005 0.819±0.009 0.720±0.015 9.552 9.448 10.621 LightGBM 0.893±0.004 0.826±0.006 0.725±0.016 8.793 8.103 8.655 ET 0.893±0.003 0.809±0.007 0.675±0.018 11.345 12.759 14.586 ExcelFormer 0.883±0.008 0.812±0.012 0.713±0.023 13.379 11.621 12.897 ResNet 0.869±0.011 0.801±0.010 0.700±0.021 14.276 14.621 13.931 MLP 0.866±0.007 0.795±0.009 0.695±0.019 16.310 15.759 15.310 TANGOS 0.865±0.010 0.797±0.009 0.698±0.019 14.931 15.931 14.310 ModernNCA 0.860±0.011 0.798±0.010 0.692±0.020 17.345 15.759 14.724 SwitchTab 0.858±0.014 0.776±0.013 0.626±0.044 18.241 19.448 19.069 FT-Transformer 0.849±0.018 0.789±0.012 0.683±0.017 18.103 17.897 16.586 T2G-Former 0.848±0.024 0.792±0.016 0.688±0.031 17.379 15.897 15.897 MLP-PLR 0.848±0.006 0.786±0.007 0.650±0.020 18.414 18.828 19.931 DCN-v2 0.846±0.021 0.771±0.020 0.633±0.058 21.793 23.000 23.103 DANets 0.844±0.039 0.770±0.040 0.631±0.077 18.897 18.862 19.966 NODE 0.844±0.058 0.754±0.060 0.535±0.175 22.069 23.862 26.000 TabR 0.842±0.028 0.789±0.013 0.688±0.024 20.759 17.931 16.483 AutoInt 0.838±0.024 0.772±0.017 0.640±0.049 22.103 21.241 21.069 TabCaps 0.834±0.022 0.788±0.014 0.636±0.031 20.069 18.103 19.207 SNN 0.831±0.027 0.762±0.023 0.595±0.063 23.379 21.138 23.034 RealMLP 0.829±0.016 0.780±0.008 0.678±0.014 21.241 19.138 18.241 TabNet 0.825±0.043 0.768±0.027 0.631±0.059 24.759 22.586 23.793 TabTransformer 0.821±0.025 0.761±0.015 0.604±0.057 21.483 20.724 22.379 GrowNet 0.756±0.024 0.689±0.019 0.497±0.034 27.621 26.655 26.138 SAINT 0.708±0.033 0.669±0.020 0.563±0.037 22.241 21.379 20.276
@@
TabZilla Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX 0.943±0.005 0.885±0.018 0.836±0.022 4.556 4.593 5.741 AutoGluon 0.933±0.001 0.871±0.002 0.803±0.008 6.778 7.148 8.593 TabICL 0.933±0.006 0.864±0.016 0.803±0.023 8.444 9.185 10.148 TabPFN-v2 0.929±0.002 0.863±0.008 0.797±0.012 7.333 8.481 8.889 XGBoost 0.929±0.003 0.863±0.005 0.789±0.011 8.741 10.074 11.185 LightGBM 0.927±0.003 0.863±0.007 0.796±0.010 10.444 9.667 10.185 RF 0.924±0.004 0.852±0.006 0.773±0.012 10.926 11.852 12.481 CatBoost 0.922±0.006 0.848±0.011 0.780±0.020 12.185 13.296 13.185 Mitra 0.915±0.010 0.841±0.015 0.758±0.037 14.074 14.074 15.037 ExcelFormer 0.915±0.012 0.861±0.014 0.802±0.029 13.815 11.444 11.444 ET 0.912±0.004 0.837±0.008 0.745±0.014 14.407 14.926 16.222 T2G-Former 0.909±0.009 0.852±0.012 0.790±0.014 14.148 14.333 14.407 TANGOS 0.909±0.006 0.841±0.014 0.776±0.032 15.148 13.704 14.148 ResNet 0.908±0.008 0.834±0.017 0.769±0.028 16.444 15.148 14.741 ModernNCA 0.907±0.007 0.850±0.012 0.794±0.017 14.222 14.556 13.926 MLP-PLR 0.906±0.005 0.847±0.008 0.773±0.016 14.963 14.889 15.926 TabR 0.904±0.014 0.853±0.014 0.793±0.029 14.963 13.185 13.630 DCN-v2 0.904±0.008 0.844±0.013 0.781±0.018 17.556 16.519 17.111 MLP 0.903±0.013 0.825±0.021 0.747±0.042 16.148 16.556 17.000 FT-Transformer 0.903±0.023 0.842±0.027 0.769±0.058 15.926 15.926 15.370 AutoInt 0.896±0.020 0.833±0.030 0.748±0.069 16.704 16.259 16.519 RealMLP 0.892±0.012 0.846±0.009 0.781±0.018 19.370 16.407 15.741 TabCaps 0.887±0.014 0.816±0.016 0.729±0.032 22.926 20.296 20.630 DANets 0.881±0.039 0.800±0.050 0.712±0.091 20.889 19.963 20.556 SNN 0.874±0.021 0.816±0.020 0.706±0.038 20.444 18.704 19.741 NODE 0.869±0.046 0.784±0.046 0.633±0.101 23.481 23.593 25.556 SwitchTab 0.860±0.030 0.764±0.045 0.660±0.073 21.889 23.519 23.148 TabNet 0.860±0.035 0.771±0.051 0.668±0.070 26.407 25.778 25.407 GrowNet 0.829±0.019 0.732±0.017 0.651±0.032 25.444 25.037 24.037 SAINT 0.824±0.013 0.764±0.017 0.680±0.046 19.148 18.852 18.556 TabTransformer 0.814±0.020 0.759±0.024 0.659±0.051 21.333 19.852 20.778
@@ 7.2 Regression
For the quantitative evaluation of regression performance, we leverage four benchmarks, including three open-source benchmark, TALENT-REG (Liu et al., 2024), PFN-REG (Hollmann et al., 2025), and CTR23 (Fischer et al., 2023). After applying filters to exclude datasets containing more than 50,000 training samples or 10,000 features, we have 33 datasets from CTR23, 28 from PFN-REG, and 99 from TALENT-REG.
@@
For comparison, we include tree-based methods, NN-based methods, AutoML frameworks, and ICL-based models, which are basically consistent with baselines employed in classification tasks described above. TabICL and TabCaps are excluded because it cannot handle regression tasks. We add another baseline, DNNR (Nader et al., 2022), which is specifically designed for regression. The model training procedure remains aligned with that employed in the classification experiments.

### v2

@@ Results.
OpenML-cc18 Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX-16M 0.939 0.893 0.811 4.339 5.323 4.548 LimiX-2M 0.936 0.891 0.803 5.065 5.387 5.919 AutoGluon 0.932 0.885 0.790 6.500 7.113 7.613 TabPFN-v2 0.929 0.886 0.790 7.887 7.726 8.161 TabICL 0.927 0.875 0.782 8.839 10.339 10.306 XGBoost 0.929 0.879 0.775 9.661 10.839 10.790 LightGBM 0.927 0.879 0.775 10.065 10.226 10.323 CatBoost 0.926 0.870 0.770 12.177 13.516 13.419 Mitra 0.920 0.866 0.743 12.952 14.419 16.048 ET 0.922 0.861 0.721 13.774 15.887 17.774 RF 0.925 0.871 0.762 13.774 13.161 14.048 ExcelFormer 0.918 0.870 0.773 13.935 13.032 12.839 TANGOS 0.910 0.863 0.759 14.161 14.129 14.290 ResNet 0.913 0.860 0.764 14.226 14.371 13.581 MLP 0.908 0.857 0.743 14.710 14.871 15.871 T2G-Former 0.908 0.859 0.748 15.919 15.500 15.484 MLP-PLR 0.896 0.858 0.734 16.694 16.435 16.790 RealMLP 0.889 0.858 0.742 17.032 15.435 14.823 FT-Transformer 0.904 0.856 0.739 17.339 16.161 17.065 ModernNCA 0.906 0.858 0.747 17.452 16.839 16.371 TabR 0.900 0.863 0.757 17.468 14.210 13.790 AutoInt 0.889 0.835 0.717 18.984 19.032 19.129 DANets 0.900 0.840 0.715 19.000 18.435 19.435 DCN-v2 0.899 0.851 0.729 19.581 18.355 18.968 SNN 0.896 0.835 0.698 21.081 20.871 21.661 SwitchTab 0.881 0.819 0.671 22.355 21.758 22.419 TabCaps 0.875 0.848 0.692 22.548 19.484 21.097 TabTransformer 0.829 0.776 0.616 23.500 22.532 23.274 NODE 0.889 0.809 0.626 24.129 25.129 27.161 SAINT 0.512 0.492 0.441 26.129 25.532 24.210 TabNet 0.863 0.816 0.666 26.226 24.065 25.823 GrowNet 0.816 0.748 0.581 26.387 26.694 25.903
@@
Table 7: Classification results on the TabArena-CLS benchmark. The best scores are shown in bold.
TabArena-CLS Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX-16M 0.849 0.877 0.597 3.455 3.394 7.000 LimiX-2M 0.846 0.876 0.594 4.879 3.333 6.727 AutoGluon 0.844 0.870 0.574 5.818 5.364 9.091 TabICL 0.840 0.870 0.553 6.879 6.394 10.576 TabPFN-v2 0.838 0.872 0.589 7.273 4.636 8.848 LightGBM 0.841 0.868 0.574 7.455 8.333 10.545 XGBoost 0.838 0.867 0.567 8.394 7.818 10.909 RF 0.837 0.864 0.558 9.818 8.303 12.061 CatBoost 0.835 0.867 0.574 10.121 7.606 9.970 ET 0.833 0.857 0.505 10.939 11.212 16.333 Mitra 0.815 0.862 0.533 12.121 10.364 14.970 ExcelFormer 0.810 0.849 0.555 14.697 16.939 14.970 MLP 0.772 0.822 0.459 18.515 17.636 19.576 TANGOS 0.791 0.844 0.522 18.576 17.848 16.000 T2G-Former 0.779 0.822 0.482 18.667 20.061 15.848 MLP-PLR 0.781 0.836 0.460 19.030 19.091 19.667 ModernNCA 0.783 0.846 0.511 19.606 20.394 16.970 AutoInt 0.769 0.826 0.474 19.697 20.636 18.576 TabR 0.785 0.842 0.510 20.152 19.606 16.091 NODE 0.769 0.792 0.352 20.303 21.424 24.333 ResNet 0.781 0.824 0.532 20.364 21.485 16.303 FT-Transformer 0.770 0.803 0.468 20.515 21.636 18.606 TabTransformer 0.739 0.781 0.438 20.606 21.424 19.091 RealMLP 0.777 0.822 0.512 20.758 21.030 14.424 DCN-v2 0.769 0.833 0.482 21.303 19.424 18.212 SNN 0.755 0.818 0.442 22.242 22.061 22.000 TabCaps 0.742 0.837 0.471 22.333 17.970 19.667 DANets 0.749 0.776 0.453 23.152 23.848 18.333 SwitchTab 0.754 0.799 0.409 23.212 24.879 22.303 SAINT 0.694 0.739 0.437 24.545 25.061 21.515 TabNet 0.709 0.789 0.438 25.848 22.364 23.182 GrowNet 0.646 0.674 0.361 26.636 28.091 23.273
@@
(a) AUC on the TabArena-CLS benchmark.
(b) Accuracy on the TabArena-CLS benchmark.
(c) F1-score on the TabArena-CLS benchmark.
Figure 11: Critical difference diagrams on the TabArena-CLS benchmark.
@@
PFN-CLS Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX-16M 0.923 0.862 0.786 2.172 3.207 3.103 LimiX-2M 0.913 0.848 0.766 4.034 6.621 5.897 TabPFN-v2 0.910 0.845 0.756 5.586 7.172 7.138 AutoGluon 0.906 0.835 0.738 5.897 7.586 7.379 TabICL 0.903 0.832 0.742 6.862 8.552 9.034 XGBoost 0.898 0.831 0.733 9.103 7.483 8.448 Mitra 0.897 0.826 0.719 9.483 10.310 12.103 LightGBM 0.893 0.826 0.725 9.586 8.724 9.345 CatBoost 0.895 0.819 0.720 10.414 10.069 11.414 RF 0.896 0.822 0.721 11.241 10.897 12.103 ET 0.893 0.809 0.675 12.138 13.517 15.483 ExcelFormer 0.883 0.812 0.713 14.310 12.379 13.759 ResNet 0.869 0.801 0.700 15.241 15.379 14.828 TANGOS 0.865 0.797 0.698 15.897 16.759 15.207 MLP 0.866 0.795 0.695 17.276 16.517 16.138 T2G-Former 0.848 0.792 0.688 18.276 16.690 16.759 ModernNCA 0.860 0.798 0.692 18.276 16.586 15.586 FT-Transformer 0.849 0.789 0.683 19.000 18.759 17.448 SwitchTab 0.858 0.776 0.626 19.172 20.379 20.000 MLP-PLR 0.848 0.786 0.650 19.310 19.759 20.931 DANets 0.844 0.770 0.631 19.862 19.690 20.862 TabCaps 0.834 0.788 0.636 21.000 19.034 20.138 TabR 0.842 0.789 0.688 21.690 18.621 17.276 RealMLP 0.829 0.780 0.678 22.172 20.069 19.138 TabTransformer 0.821 0.761 0.604 22.379 21.517 23.310 DCN-v2 0.846 0.771 0.633 22.759 23.759 24.069 NODE 0.844 0.754 0.535 23.000 24.690 26.966 AutoInt 0.838 0.772 0.640 23.069 21.966 22.034 SAINT 0.708 0.669 0.563 23.207 22.138 21.207 SNN 0.831 0.762 0.595 24.345 21.931 23.966 TabNet 0.825 0.768 0.631 25.759 23.552 24.759 GrowNet 0.756 0.689 0.497 28.621 27.586 27.138
@@
TabZilla Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX-16M 0.943 0.885 0.836 4.815 4.963 6.185 LimiX-2M 0.938 0.883 0.832 5.667 6.222 6.630 AutoGluon 0.933 0.871 0.803 7.370 7.556 9.148 TabPFN-v2 0.929 0.863 0.797 8.000 9.037 9.481 TabICL 0.933 0.864 0.803 9.111 9.815 10.815 XGBoost 0.929 0.863 0.789 9.407 10.667 11.926 LightGBM 0.927 0.863 0.796 11.185 10.296 10.852 RF 0.924 0.852 0.773 11.704 12.593 13.185 CatBoost 0.922 0.848 0.780 12.926 13.926 13.926 ExcelFormer 0.915 0.861 0.802 14.667 12.259 12.296 Mitra 0.915 0.841 0.758 14.926 14.815 15.778 T2G-Former 0.909 0.852 0.790 14.926 14.852 15.148 ModernNCA 0.907 0.850 0.794 14.963 15.333 14.630 ET 0.912 0.837 0.745 15.333 15.741 17.037 TabR 0.904 0.853 0.793 15.741 13.963 14.370 MLP-PLR 0.906 0.847 0.773 15.741 15.519 16.704 TANGOS 0.909 0.841 0.776 15.926 14.444 14.963 FT-Transformer 0.903 0.842 0.769 16.741 16.444 16.074 MLP 0.903 0.825 0.747 17.000 17.333 17.852 ResNet 0.908 0.834 0.769 17.333 16.000 15.556 AutoInt 0.896 0.833 0.748 17.593 16.852 17.370 DCN-v2 0.904 0.844 0.781 18.333 17.370 17.889 SAINT 0.824 0.764 0.680 20.074 19.667 19.370 RealMLP 0.892 0.846 0.781 20.259 17.259 16.556 SNN 0.874 0.816 0.706 21.333 19.481 20.667 DANets 0.881 0.800 0.712 21.815 20.852 21.444 TabTransformer 0.814 0.759 0.659 22.222 20.630 21.593 SwitchTab 0.860 0.764 0.660 22.815 24.444 24.111 TabCaps 0.887 0.816 0.729 23.926 21.222 21.556 NODE 0.869 0.784 0.633 24.444 24.481 26.519 GrowNet 0.829 0.732 0.651 26.407 26.000 25.037 TabNet 0.860 0.771 0.668 27.370 26.778 26.407
@@ 7.2 Regression
For the quantitative evaluation of regression performance, we leverage four benchmarks, including three open-source benchmark, TALENT-REG (liu2024talenttabularanalyticslearning), PFN-REG (hollmann2025tabpfn_nature),TabArena-REG (erickson2025tabarena), and CTR23 (fischer2023openml_ctr23). After applying filters to exclude datasets containing more than 50,000 training samples or 10,000 features, we have 33 datasets from CTR23, 28 from PFN-REG, 13 from TabArena-REG, and 99 from TALENT-REG.
@@
For comparison, we include tree-based methods, NN-based methods, AutoML frameworks, and ICL-based models, which are basically consistent with baselines employed in classification tasks described above. TabICL and TabCaps are excluded because it cannot handle regression tasks. We add another baseline, DNNR (nader2022dnnr), which is specifically designed for regression. The model training procedure remains aligned with that employed in the classification experiments.